# In Search of Trees: Decision-Tree Policy Synthesis for Black-Box Systems via Search

### v1

@@ Abstract
Decision trees, owing to their interpretability, are attractive as control policies for (dynamical) systems. Unfortunately, constructing, or synthesising, such policies is a challenging task. Previous approaches do so by imitating a neural-network policy, approximating a tabular policy obtained via formal synthesis, employing reinforcement learning, or modelling the problem as a mixed-integer linear program. However, these works may require access to a hard-to-obtain accurate policy or a formal model of the environment (within reach of formal synthesis), and may not provide guarantees on the quality or size of the final tree policy. In contrast, we present an approach to synthesise optimal decision-tree policies given a black-box environment and specification, and a discretisation of the tree predicates, where optimality is defined with respect to the number of steps to achieve the goal. Our approach is a specialised search algorithm which systematically explores the (exponentially large) space of decision trees under the given discretisation. The key component is a novel pruning mechanism that significantly reduces the search space. Our approach represents a conceptually novel way of synthesising small decision-tree policies with optimality guarantees even for black-box environments with black-box specifications.
@@ 1 Introduction
In contrast, we consider a unique setting: derive 1) small decision-tree policies when 2) the model and specification of the environment is black-box whilst 3) providing optimality guarantees of performance. Every work we are aware of will violate at least one of these three points.
We note that our work is applicable to deterministic black-box systems. Such systems pose a challenging controllability problem, as providing exact guarantees implies searching in an exponentially large space.
@@
A key component is our novel trace-based pruning mechanism that discards a large portion of the search space by runtime analysis. This is made possible by exploiting the decision-tree structure and considering the execution of the tree policy: even though the environment is black-box, examining the trace allows us to understand how the decision tree is used, and discard trees that are guaranteed to not lead to a better trace. This allows us to reduce the search space without sacrificing optimality even though our model and specification are given as a black box. In the previous example, depending on the concrete trace, our trace-based pruning might be able to determine that it is possible to discard the third tree from consideration without missing a better tree only by observing the trace produced by the first and second tree. In practice this can lead to order-of-magnitude reductions in runtime.
We implemented our approach and evaluate it on classical control benchmarks. The experiments demonstrate significant reductions obtained with our trace-based pruning, and illustrate that small and optimal decision trees may be constructed within reasonable time. We further analyse the scalability of our algorithm in terms of the number of predicates (granularity) and the size of the tree, both of which have an exponential influence on the runtime. Nevertheless, the experiments show the runtime is within practical use.
@@ 2 Related Work
Our work covers a unique setting: constructing decision trees given a deterministic black-box system whilst providing optimal performance guarantees. As such, there are no directly applicable works that we are aware of in this setting. Nevertheless, to illustrate the challenges of our setting, we discuss previous works for synthesising decision-tree policies, albeit not fitting into our setting.
@@ 3 Preliminaries
A state $S=(s_{1},s_{2},\dots,s_{d})^{\top}\in\mathcal{S}$ is a $d$-dimensional real-valued vector from a bounded state space $\mathcal{S}\subseteq\mathbb{R}^{d}$, where each state dimension $s_{i}$ belongs to an interval $s_{i}\in[\ell_{i},u_{i}]$ of a lower and an upper bound. An action $a\in A$ comes from a finite set $A\subseteq\mathbb{Z}$ of integer-valued actions.
An environment is a function $E\colon\mathcal{S}\times A\to\mathcal{S}$ that takes as input a state $S$ and an action $a$ and computes a trajectory until asked to output the next observable state $S^{\prime}=E(S,a)$. In our setting, the environment is treated as a black box, i.e., we are agnostic to its dynamics.
A policy is a function $\pi\colon\mathcal{S}\to A$ that chooses an action based on an input state. We write $\mathcal{C}_{A}$ for the set of all policies over action set $A$. In this work, we are concerned with the special case of a decision-tree policy, which is given in the form of a binary tree where each inner node is called a predicate node and each leaf node is called an action node. Each predicate node is associated with a function $\mathcal{S}\to\mathbb{B}$, where $\mathbb{B}$ is the Boolean domain. Each action node is associated with one of the available actions $a\in A$. Given a state $S$ and a node of the tree, a decision-tree policy $\pi$ computes the action $a=\pi(S)$ using the following recursive procedure, starting at the root node. If the current node is an action node, the associated action is returned. Otherwise, the current node is a predicate node. If the state $S$ makes the predicate evaluate to true (false), the procedure continues with the left (right) child node.
Given an environment $E$, a decision-tree policy $\pi$, an initial state $S_{0}$, and a bound $k\in\mathbb{N}$, the trace with $k$ steps is the sequence of observed states $\tau=S_{0},S_{1},S_{2},\dots,S_{k}$, obtained by applying the sequence of actions given by $\pi$: $S_{i}=E(S_{i-1},\pi(S_{i-1}))$, for $i=1,\ldots,k$. We write $\mathcal{T}$ for the set of all traces.
@@ 4 Optimal Decision-Tree Policies
To simplify the exposition, we focus our discussion on the case of a single initial state $S_{0}$. We generalise the problem to multiple initial states in Section 5.4.
@@ 4.1 Discretised Decision-Tree Predicates
The space of all decision trees is infinitely large. By discretising the tree predicates, we obtain a finite (but exponentially large) space. We restrict our attention to (axis-aligned) predicates of the form $[s_{i}\geq v_{0}+m\cdot v_{+}]$, where $s_{i}$ is the i-th state dimension, $v_{0},v_{+}\in\mathbb{R}$ are real-valued constants, and $m\in\mathbb{N}$ is a positive integer. Since our state space is bounded, we obtain a finite number of non-equivalent predicates. We write $\mathcal{P}$ for the set of all (tree) predicates.
@@
The increment value controls the discretisation resolution: a smaller value yields more predicates, which increases the space of considered decision trees but potentially allows to find better trees. The algorithm’s runtime is sensitive to the larger search space, and hence a practical balance is needed. Experimentally we show that our approach can handle reasonably small increments.
@@ 4.2 Specification
Given environment $E$, decision-tree policy $\pi$, and initial state $S_{0}$, we consider a specification to determine whether $\pi$ satisfies the specification. We assume that the specification is given in terms of traces, again in the form of a black-box specification function $P\colon\mathcal{T}\to\mathbb{B}$. In order to effectively determine whether $\pi$ satisfies the specification, we restrict the class of specifications we consider. A bounded-time specification with a bound $k\in\mathbb{N}$ has the property that every trace $\tau$ of length $k$ either satisfies or violates the specification. As a consequence, we are guaranteed to obtain a Boolean verdict from a trace of length at most $k$. We call the (unique) trace of length $k$ the witness trace.
@@ 4.3 Optimality
So far, we were only interested in finding any policy that satisfies a given specification. In general, there may exist multiple solutions. We are interested in identifying an optimal policy. For that, we assume a fitness function, which is a partial order ${\succeq}\colon\mathcal{T}\times\mathcal{T}$ to compare two traces. A trace that satisfies the specification always precedes a trace that violates the specification. The fitness function induces another partial order ${\succeq}\colon\mathcal{C}_{A}\times\mathcal{C}_{A}$ to compare two policies, as follows. We say that $\pi_{1}$ is strictly better than $\pi_{2}$, written $\pi_{1}\succ\pi_{2}$, if one of the following conditions holds: 1) The witness $\tau_{1}$ has strictly better fitness than the witness $\tau_{2}$. 2) Both witnesses have the same fitness, and $\pi_{1}$ is strictly smaller.
@@ 5 Synthesis of Optimal Decision-Tree Policies
For computing an optimal decision-tree policy that solves Problem 1, a naive procedure is to enumerate all possible decision trees and evaluate them. By fixing an upper limit of the number of tree nodes, this procedure terminates. However, the number of available trees is exponential, rendering this procedure infeasible. Our main contribution is an efficient instantiation of this procedure.
@@ 5.1 Searching In the Space of Decision Trees
Our algorithm to enumerate decision trees is based on backtracking search. We represent the search space using backtracking variables $b_{i}$, where each variable is associated with a node in the tree. The possible values that can be assigned to a backtracking variable depend on the type of node which the variable is associated with: predicate nodes may be assigned a predicate from the set of discretised predicates, whereas action nodes may be assigned an action from the set of available actions.
@@
Assigning all backtracking variables thus results in a particular decision-tree policy, and consequently, enumerating all possible assignments to the backtracking variables corresponds to all possible policies in our available space. When enumerating a policy, it is used in combination with the black-box environment to compute the witness trace and subsequently evaluate the quality of the policy. Finally, the best policy is returned as the result.
For a tree with a fixed shape and $n$ predicate nodes, $|A|$ number of actions, and $|\mathcal{P}|$ number of discretised predicates, the size of the total search space is $\mathcal{O}(|\mathcal{P}|^{n}\cdot|A|^{n+1})$. Our key contributions are techniques for reducing this exponentially large search space in practice.
@@ 5.3 Trace-Based Pruning
Given a backtracking variable $b_{i}$ associated with a predicate node, the algorithm needs to select the next predicate to assign to the variable. A naive approach would be to simply select the next bigger one, e.g., from the example in Section 5.2, assign predicate $[s_{1}\geq 2]$ after considering $[s_{1}\geq 1]$. However, we can leverage information about previous traces to avoid explicitly considering predicates that are guaranteed to not result in a trace that has not been previously observed.
@@
To reiterate, for each backtracking variable $b_{i}$ associated with a predicate node with current predicate $[s_{j}\geq v]$, we store a value $d_{i}\in\mathbb{R}$, which tracks the minimum value of state dimension $s_{j}$ for which the predicate was evaluated to true amongst all traces that were considered since the predicate $[s_{j}\geq v]$ has been assigned. Note that selecting a new threshold value for the predicate that is smaller than $d_{i}$ is guaranteed to result in a trace already observed. Note that it is not necessary to track the values where the predicate evaluates to false, since the predicates are explored in increasing order of the thresholds and as such the future predicates would also evaluate to false on those values.
Initially, the distance value $d_{i}$ is set to undefined each time a new predicate $[s_{j}\geq v]$ is assigned as part of the search. The first time the node observes a state where its predicate is satisfied in a trace, the distance value $d_{i}$ is set to the corresponding value of $s_{j}$. Each subsequent time the predicate is satisfied, $d_{i}$ is updated to the smallest value for which the predicate still evaluates to true.
@@
18:        $i$leftarrow bvar.index
19:        bvar.distance $\leftarrow\min(\pi_{local}$.nodes[$i$].distance, bvar.distance))))
@@ 5.4 Extensions
As a result, due to the interaction with pruning, finding the optimal tree with respect to multiple initial states may lead to lower running times than with a single initial state.
@@ 6 Experimental Study
We consider three classical control problems: CartPole, MountainCar, and pendulum. The environment behaviour is defined as in Gymnasium [^1] https://github.com/Farama-Foundation/Gymnasium. MountainCar and pendulum are minimisation problems, whereas CartPole is a maximisation problem. We set the maximum trace size to $10{,}000$, which is mainly only relevant for CartPole since it maximises the trace size, while for the other two environments, most traces are cut well before the limit due to early stopping. The control actions are limited to two choices, e.g., apply maximum force in one or the other direction. Our implementation only queries the environment in a black-box fashion. We generated the initial states randomly within a specified range; see Appendix A for details about the parameters, environment description, and a sample of trees produced by our approach.
To reiterate, as discussed in Section 2, we work with a unique setting where we 1) synthesise decision-tree policies, 2) only require black-box access to the environment and the specification, and 3) provide guarantees on performance under the tree definition, e.g., the policy that minimises the time taken, or prove that no such policy exists. Every work we are aware of violates at least one of these points. Consequently, while direct comparisons with other works may be done by relaxing the requirements of our setting, this brings considerable caveats that result in comparisons that we argue are not meaningful with respect to our contribution. For these reasons, we focus on demonstrating the feasibility of our approach and its scalability.
Our code base is written in (pure) Rust 1.77.0. The experiments were run on consumer-grade hardware (Intel(R) Xeon(R) W-10855M @ 2.8 GHz). Our code will be made publicly available in due course.
@@ 6.1 Experiment #1: Trace-Based Pruning
Table 1: Summary of trace-based pruning experiments. Our pruning leads to a substantial runtime reduction.
@@
Table 2: Summary of experiments varying the granularity of predicates. Runtime increases with finer discretisation, with the exception of CartPole, where a finer discretisation leads to finding a perfect tree faster.
Runtime (seconds) No. of trees
Environment 5 10 15 20 5 10 15 20
@@
Table 3: Summary of experiments varying the number of predicate nodes. Runtime and performance (not depicted) increase.
Runtime (seconds) No. of trees
Environment 3 4 5 6 3 4 5 6

### v2

@@ Abstract
Decision trees, owing to their interpretability, are attractive as control policies for (dynamical) systems. Unfortunately, constructing, or synthesising, such policies is a challenging task. Previous approaches do so by imitating a neural-network policy, approximating a tabular policy obtained via formal synthesis, employing reinforcement learning, or modelling the problem as a mixed-integer linear program. However, these works may require access to a hard-to-obtain accurate policy or a formal model of the environment (within reach of formal synthesis), and may not provide guarantees on the quality or size of the final tree policy. In contrast, we present an approach to synthesise optimal decision-tree policies given a deterministic black-box environment and specification, a discretisation of the tree predicates, and an initial set of states, where optimality is defined with respect to the number of steps to achieve the goal. Our approach is a specialised search algorithm which systematically explores the (exponentially large) space of decision trees under the given discretisation. The key component is a novel trace-based pruning mechanism that significantly reduces the search space. Our approach represents a conceptually novel way of synthesising small decision-tree policies with optimality guarantees even for black-box environments with black-box specifications.
Codehttps://doi.org/10.5281/zenodo.14601859
@@ 1 Introduction
In contrast, we consider a unique setting: derive 1) small decision-tree policies when 2) the model and specification of the environment is a deterministic black-box whilst 3) providing optimality guarantees of performance, under the discretisation of tree predicates and initial states. Every work we are aware of will violate at least one of these three points.
Our work is for deterministic black-box systems, which pose a challenging controllability problem, as exact guarantees imply searching in an exponentially large space.
@@
A key component is our novel trace-based pruning mechanism that discards a large portion of the search space by runtime analysis. It exploits the decision-tree structure by considering the execution of the tree policy: even though the environment is black-box, examining the trace allows us to understand how the decision tree is used, and discard trees that are guaranteed to not lead to a better trace. This allows us to reduce the search space without sacrificing optimality even though our model and specification are given as a black box. In the previous example, depending on the concrete trace, our trace-based pruning might be able to determine that it is possible to discard the third tree from consideration without missing a better tree only by observing the trace produced by the first and second tree. In practice this can lead to order-of-magnitude reductions in runtime.
We implemented our approach and evaluate it on classical control benchmarks. The experiments demonstrate significant reductions obtained with our trace-based pruning, and illustrate that small and optimal decision trees may be constructed within reasonable time. We further analyse the scalability of our algorithm in terms of the number of predicates (granularity) and the size of the tree, both of which have an exponential influence on the runtime. The experiments show the runtime is within practical use.
@@ 2 Related Work
Our work covers a unique setting: constructing decision trees given a deterministic black-box system whilst providing optimal performance guarantees. There are no directly applicable works that we are aware of in this setting. To illustrate the challenges, we discuss works for synthesising decision-tree policies, albeit not fitting into our setting.
@@ 3 Preliminaries
A state $S=(s_{1},s_{2},\dots,s_{d})^{\top}\in\mathcal{S}$ is a $d$-dimensional real-valued vector from a bounded state space $\mathcal{S}\subseteq\mathbb{R}^{d}$, where each state dimension $s_{i}$ belongs to an interval $s_{i}\in[\ell_{i},u_{i}]$. An action $a\in A$ comes from a finite set $A\subseteq\mathbb{Z}$ of integer-valued actions. An environment is a function $E\colon\mathcal{S}\times A\to\mathcal{S}$ that takes as input a state $S$ and an action $a$ and computes a trajectory until asked to output the next observable state $S^{\prime}=E(S,a)$. In our setting, the environment is treated as a black box, i.e., we are agnostic to its dynamics.
A policy is a function $\pi\colon\mathcal{S}\to A$ that chooses an action based on an input state. We write $\mathcal{C}_{A}$ for the set of all policies over action set $A$. We are concerned with the special case of a decision-tree policy, which is given in the form of a binary tree where each inner node is called a predicate node and each leaf node is called an action node. Each predicate node is associated with a function $\mathcal{S}\to\mathbb{B}$, where $\mathbb{B}$ is the Boolean domain. Each action node is associated with one of the available actions $a\in A$. Given a state $S$ and a node of the tree, a decision-tree policy $\pi$ computes the action $a=\pi(S)$ using the following recursive procedure, starting at the root node. If the current node is an action node, the associated action is returned. Otherwise, the current node is a predicate node. If the state $S$ makes the predicate evaluate to true (false), the procedure continues with the left (right) child node.
Given an environment $E$, a decision-tree policy $\pi$, an initial state $S_{0}$, and a bound $k\in\mathbb{N}$, the trace with $k$ steps is the sequence of observed states $\tau=S_{0},S_{1},\dots,S_{k}$, obtained by applying the sequence of actions given by $\pi$: $S_{i}=E(S_{i-1},\pi(S_{i-1}))$, for $i=1,\ldots,k$. We write $\mathcal{T}$ for the set of all traces.
@@ 4 Optimal Decision-Tree Policies
We focus on the case of a single initial state $S_{0}$, and generalise the problem to multiple initial states in Section 5.4.
@@ 4.1 Discretised Decision-Tree Predicates
The space of all decision trees is infinitely large. By discretising the tree predicates, we obtain a finite (but exponentially large) space. We restrict our attention to (axis-aligned) predicates of the form $[s_{i}\geq v_{0}+m\cdot v_{+}]$, where $s_{i}$ is the i-th state dimension, $v_{0},v_{+}\in\mathbb{R}$ are real-valued constants, and $m\in\mathbb{N}$ is a positive integer. Since our state space is bounded, we obtain a finite number of nonequivalent predicates. We write $\mathcal{P}$ for the set of all (tree) predicates.
@@
Considering a finer discretisation yields more predicates, which increases the space of considered decision trees, but allows potentially finding better trees. The algorithm’s runtime is sensitive to the size of the search space, and hence a practical balance is needed. Experimentally we show that our approach can handle reasonably small increments.
@@ 4.2 Specification
Given a deterministic environment $E$, decision-tree policy $\pi$, and initial state $S_{0}$, we consider a specification to determine whether $\pi$ satisfies the specification. We assume that the specification is given in terms of traces, again in the form of a black-box specification function $P\colon\mathcal{T}\to\mathbb{B}$. In order to effectively determine whether $\pi$ satisfies the specification, we restrict the class of specifications we consider. A bounded-time prefix-closed specification with a bound $k\in\mathbb{N}$ has the property that every trace $\tau$ of length $k$ either satisfies or violates the specification (“bounded-time”), and whenever a trace $\tau$ yields either of these verdicts, then any longer trace with the prefix $\tau$ yields the same verdict (“prefix-closed”). As a consequence, we are guaranteed to obtain a Boolean verdict from a trace of length at most $k$. We call the (unique) trace of length $k$ the witness trace.
@@ 4.3 Optimality
So far, we were only interested in finding any policy that satisfies a given specification. In general, there may exist multiple solutions. We are interested in identifying an optimal policy. For that, we assume a fitness function, which is a partial order ${\succeq}\colon\mathcal{T}\times\mathcal{T}$ to compare two traces. A trace that satisfies the specification always precedes a trace that violates the specification. The fitness function induces another partial order ${\succeq}\colon\mathcal{C}_{A}\times\mathcal{C}_{A}$ to compare two policies. We say that $\pi_{1}$ is strictly better than $\pi_{2}$, written $\pi_{1}\succ\pi_{2}$, if one of the following conditions holds: 1) The witness $\tau_{1}$ has strictly better fitness than the witness $\tau_{2}$ (i.e., $\tau_{1}\succ\tau_{2}$). 2) Both witnesses have the same fitness, and $\pi_{1}$ is a strictly smaller tree.
@@ 5 Synthesis of Optimal Decision-Tree Policies
For computing an optimal decision-tree policy that solves Problem 1, a naive procedure is to enumerate all possible decision trees and evaluate them. By fixing an upper limit of the number of nodes and considering a discrete set of predicates, this procedure terminates. However, the number of trees is exponential, rendering this procedure infeasible. Our contribution is an efficient instantiation of this procedure.
@@ 5.1 Searching In the Space of Decision Trees
Our algorithm to enumerate decision trees is based on backtracking search. We represent the search space using backtracking variables $b_{i}$, where each variable is associated with a node in the tree. The possible values that can be assigned to a variable depend on the node type which the variable is associated with: predicate nodes may be assigned a predicate from the set of discretised predicates, whereas action nodes may be assigned an action from the set of available actions.
@@
Assigning all backtracking variables results in a decision-tree policy. Consequently, enumerating all possible assignments to the backtracking variables corresponds to all possible policies in our available space. When enumerating a policy, it is used in combination with the black-box environment to compute the witness trace and evaluate the quality of the policy. Finally, the best policy is returned as the result.
For a tree with a fixed shape and $n$ predicate nodes, $|A|$ number of actions, and $|\mathcal{P}|$ number of discretised predicates, the size of the total search space is $\mathcal{O}(|\mathcal{P}|^{n}\cdot|A|^{n+1})$. Our contribution reduces this large search space in practice.
@@ 5.3 Trace-Based Pruning
Given a backtracking variable $b_{i}$ associated with a predicate node, the algorithm needs to select the next predicate to assign to the variable. A naive approach would be to simply select the next larger threshold, e.g., from the example in Section 5.2, assign predicate $[s_{1}\geq 2]$ after considering $[s_{1}\geq 1]$. However, we can leverage information about previous traces to avoid explicitly considering predicates that are guaranteed to not result in a trace that has not been previously observed.
@@
To reiterate, for each backtracking variable $b_{i}$ associated with a predicate node with current predicate $[s_{j}\geq v]$, we store a value $d_{i}\in\mathbb{R}$, which tracks the minimum value of state dimension $s_{j}$ for which the predicate was evaluated to true amongst all traces that were considered since the predicate $[s_{j}\geq v]$ has been assigned. Note that selecting a new threshold value for the predicate that is smaller or equal than $d_{i}$ is guaranteed to result in traces already observed. Note that it is not necessary to track the values where the predicate evaluates to false, since the predicates are explored in increasing order of the thresholds and as such the future predicates would also evaluate to false on those values.
Initially, the distance value $d_{i}$ is set to undefined each time a new predicate $[s_{j}\geq v]$ is assigned as part of the search. The first time the node observes a state where its predicate is satisfied in a trace, the distance value $d_{i}$ is set to the corresponding value of state dimension $s_{j}$. Each subsequent time the predicate is satisfied, $d_{i}$ is updated to the smallest value for which the predicate still evaluates to true.
@@
18:        n $\leftarrow\pi_{local}$.nodes[bvar.index]
19:        bvar.distance $\leftarrow\min($n.distance, bvar.distance))))
@@ 5.4 Extensions
As a result, due to the interaction with pruning, finding an optimal tree with respect to multiple initial states may lead to sub-linear runtime with respect to the number of states.
@@ 6 Experimental Study
We consider three classical control problems: CartPole, MountainCar, and pendulum. The environment behaviour is defined as in Gymnasium [^1] https://github.com/Farama-Foundation/Gymnasium, with the adjustment that we maximise (CartPole - at most 10k steps) or minimise (MountainCar and pendulum) the trace length rather than use their reward functions. The control actions are limited to two choices, e.g., apply maximum force in one or the other direction. Our implementation only queries the environment in a black-box fashion. We generated the initial states randomly within a specified range; see the Appendix for details about the parameters, environment description, and a sample of trees produced by our approach.
To reiterate, as discussed in Section 2, we work with a unique setting where we 1) synthesise decision-tree policies, 2) only require black-box access to the deterministic system, and 3) provide guarantees on performance under the tree definition, e.g., the policy that minimises the time taken, or prove that no such policy exists. Every work we are aware of violates at least one of these points. Consequently, while direct comparisons with other works may be done by relaxing the requirements of our setting, this brings considerable caveats that result in comparisons that we argue are not meaningful with respect to our contribution. For these reasons, we focus on demonstrating the feasibility of our approach and its scalability.
We implemented our approach ‘Broccoli’ in Rust 1.77.0. The experiments were run on consumer-grade hardware (Xeon(R) W-10855M @ 2.8 GHz). The code is public.
@@ 6.1 Experiment #1: Trace-Based Pruning
Table 1: Runtime and variance with and without trace-based pruning. Pruning substantially reduces runtime. Multiple initial states demonstrate only a sub-linear runtime increase.
@@
Table 2: Runtime increases as we increase the number of predicates per state dimension (X), with the exception of CartPole, where finer discretisation sometimes finds a perfect tree faster. Variance remains similar as in Table 1 proportional to runtime.
Runtime (seconds) and variance No. of trees
Environment X=5 X=10 X=15 X=20 X=5 X=10 X=15 X=20
@@
Table 3: Runtime and performance (not depicted) increase as we increase the number of nodes (N).
Runtime (seconds) and variance No. of trees
Environment N=3 N=4 N=5 N=6 N=3 N=4 N=5 N=6